OpenStudy (anonymous):

@Mathematics What is the range of f(x) = -x2 + 4 if the domain is {2,0,1}? A){0,4,3} B){0,4,5} C){8,4,3} D){8,4,5}

OpenStudy (ash2326):

i'm assuming ur ques is f(x)=-x^2+4 put x=2 you get f(x)=-2^2+2=0 put x=0 f(x)=4 put x=1 f(x)=-1+4=3 so range is 0,4,3 option a

OpenStudy (mathteacher1729):

Make a little table: f(x) = -x^2 + 4 and x = 2, 0 , 1 Now find f(2) , f(0), f(1). "f evaluated at x = 2" , "f evaluated at x = 0" etc.